OpenStudy (xoxolove456):

A plane travels 210 mph in still air. It takes 2 hours longer to travel 425 miles flying against the wind then it does flying with the wind. Find the speed of the wind.

OpenStudy (xoxolove456):

Round to nearest tenth

OpenStudy (dumbcow):

\[t = \frac{d}{r}\] \[\frac{425}{210-w} - \frac{425}{210+w} = 2\]

OpenStudy (xoxolove456):

So it's just 2?

OpenStudy (dumbcow):

??? No you have to algebraically solve for "w", which represents speed of wind